OpenStudy (anonymous):

solve the system of equations x+y=7 x-y=3 the answer should be in an ordered pair

OpenStudy (anonymous):

You would solve this by Elimination. Which means that

OpenStudy (anonymous):

Isolate for x in the second equation: x= 3+y. Insert 3+y into the first equation to solve for y: 3+y + y = 7 2y=4 y=2 x=3+2y = 3+ 4 = 7.

OpenStudy (anonymous):

Dang it :P lol

OpenStudy (anonymous):

x+y=7 time (-1) then -x-y=-7 -x-y=-7 ADD x-y=3 ----------------- -2y = -4 y= -4/-2 y=2 subtitution y =2 slove for x x+2=7 x= 7-2 x =5

OpenStudy (toxicsugar22):

x=5,y=-2

OpenStudy (anonymous):

x=5 y=2

OpenStudy (anonymous):

I sure it correct

OpenStudy (anonymous):

how would you graph this?

OpenStudy (toxicsugar22):

yeah ur right it is corect

OpenStudy (anonymous):

The x is the slope and the y is the y intercept (:

OpenStudy (anonymous):

the rest of my question wants to know if this is dependent or independent and alsoconsistent or inconsistent

OpenStudy (toxicsugar22):

it is dependent and it is consistent

OpenStudy (anonymous):

CRAP!! i forgot the definition.... >.<

OpenStudy (anonymous):

can anyone graph it and send me a link so i can see how it should look?

OpenStudy (anonymous):

sorry I wrong page

OpenStudy (anonymous):

u have grid paper?

OpenStudy (anonymous):

no i do it on the computer

OpenStudy (anonymous):

u need rewrite to slope intercept than draw grap on the grid paper,

OpenStudy (anonymous):

Yeah, there was a slight mistake in that I wrote x=3+2y instead of x=3+y. Honestly, graphing a function like this shouldn't be any problem at all. y=x-3. Start at y(0)=-3, y(1)= -2, y(2)=-1, y(3)=0, and so on... Isolate for y, (y=x-3), and insert whatever x you need in order to get the information you need.

OpenStudy (anonymous):

x+y=7 , y= -x +7 x-y=3 , y = x-3

OpenStudy (anonymous):

let me send u grap , u will see intercept at (5,2)

OpenStudy (anonymous):

sorry y=2 on the grap I write y =3

OpenStudy (anonymous):

Isn't it a bit more productive to show this person the way to the anser instead of spoonfeeding him?

OpenStudy (anonymous):

Answer.

OpenStudy (anonymous):

Answer?